OpenStudy (lgbasallote):

If \(\vec A\) = 5, \(\vec B\) = 8, and \(\alpha\) = 30\(^o\), determine the scalar product of the two vectors shown

OpenStudy (lgbasallote):



OpenStudy (lgbasallote):

do i just use \[5 \times 8 \times \cos (30^o)\]

OpenStudy (lgbasallote):

no...i don't think that works....

OpenStudy (anonymous):

A.B=ABCOSx A.B=5*8cos30 A.B=40*0.866 A.B=34.64

OpenStudy (lgbasallote):

i think it should be \[5 \times 8 \times \cos (150^o)\]

OpenStudy (anonymous):

@lgbasallote lets suppose this two A and B vectors are the force pulling a body from two sides and the scalar product means finding teh work done...

OpenStudy (lgbasallote):

ugh please don't relate it to physics =_=

OpenStudy (anonymous):